# Adaptive Depth Networks with Skippable Sub-Paths

### v2

@@ 4.3 On-Device Performance
Figure 5(b) shows the performance in Nvidia Jetson Orin Nano. The inference latency and energy consumption of ResNet50-ADN is compared to S-ResNet50, a representative width-adaptation network. The result shows that depth-adaptation of ResNet50-ADN is highly effective in accelerating inference speeds and reducing energy consumption. For example, in ResNet50-ADN, reducing FLOPs by 38% through depth adaptation reduces both inference latency and energy consumption by 35%. In contrast, even though S-ResNet50 can reduce FLOPs by up to 93% by adjusting its width, it only achieves up to 9% acceleration in practice.
@@ 4.4 Ablation Study
Table 2: Ablation analysis with ResNet50-ADN and ViT-b/32-ADN. Applied components are checked. $\downarrow$ and $\uparrow$ in parentheses are comparisons to non-adaptive individual networks.
self- skip-aware ResNet50 Acc@1 (%) ViT-b/32 Acc@1 (%)
@@
O 76.1% ($\downarrow$ 0.6%) 74.9 % ($\downarrow$ 0.1%) 76.4% ($\uparrow$0.5%) 74.3% ($\uparrow$0.5%)
O 76.6% ($\downarrow$ 0.1%) 75.1% ($\uparrow$ 0.1%) 76.0% ($\uparrow$0.1%) 74.3% ($\uparrow$0.5%)
O O 77.6% ($\uparrow$ 0.9%) 76.1% ($\uparrow$ 1.1%) 76.6% ($\uparrow$ 0.8%) 74.3% ($\uparrow$0.5%)
We first investigate the influence of two key components of the proposed adaptive depth networks: (1) self-distillation of sub-paths and (2) skip-aware BNs/LNs. When our self-distillation method is not applied, the loss of the base-net, or $loss_{base}$, in Algorithm 1 is modified to $criterion(\mathbf{y},\mathbf{\hat{y}}_{base})$. Table 2 shows the results. For ResNet50, when neither of them is applied, the inference accuracy of the super-net and the base-net is significantly lower than non-adaptive individual networks by 1.5% and 2.8%, respectively. This result shows the difficulty of joint training sub-networks for adaptive networks. When one of the two components is applied individually, the performance is still slightly worse than individual networks’. In the last row, when both self-distillation and skip-aware BNs are applied together, ResNet50-ADN achieves significantly better performance than individual networks, both in the super-net and the base-net. The results with ViT-b/32 show that switchable layer normalization has similar effect in vision transformers.
@@
Self-Distillation Strategies: Our self-distillation approach in Algorithm 1 exploits only two sub-networks exclusively as a teacher and a student. Specifically, the super-nets, or FFFF, acts as the the teacher and the base-nets, or TTTT, becomes the student. The purpose of this strategy is not to train only those two sub-networks, but rather to train skippable sub-paths in a way that minimally modifies the feature distribution, as demonstrated in Figure 3. To investigate the effect of self-distillation strategies, we conduct an experiment in Table 3. In every training iteration, rather than exclusively using FFFF and TTTT sub-networks for self-distillation, we randomly sample either a teacher, a student, or both from $2^{4}$ sub-networks. These randomly sampled sub-networks are trained explicitly through self-distillation. However, as shown in Table 3, all sub-networks, both from ResNet50-ADN and ViT/b-32-ADN, perform significantly better when our self-distillation strategy is applied for training. Even though some sub-networks (such as TTTF, TTFF, and TTTF) are not trained explicitly with our self-distillation strategy, they still outperform their counterpart sub-networks trained through random sampling. This result demonstrates that our method of training sub-paths is more effective than training target sub-networks themselves.
Figure 6: Pareto frontiers of ResNet50-ADNs trained with varying ratios between mandatory and skippable sub-paths. Total number of blocks remains unchanged.
Lengths of Mandatory Sub-Paths: In Figure 6, we investigate the effect of varying the length ratio between the mandatory and the skippable sub-paths. (Details are in Table 6.) If mandatory sub-paths become shorter than skippable sub-paths, smaller sub-networks can be selected since more layers can be skipped. However, the result in Figure 6 shows that this configuration significantly degrades performance of all sub-networks. Since every sub-network shares parameters of mandatory sub-paths, low inference capability of shallow mandatory sub-paths affects all sub-networks. This implies that maintaining certain depths in mandatory sub-paths is essential for effective inference. In contrast, increasing the length of mandatory sub-paths further does not yield further performance improvement and only reduce the range of FLOPs adaptation.
@@ 5 Conclusions
We propose a practical approach to adaptive depth networks that can be applied to both CNNs and transformers with minimal training effort. We provide a general principle and formal explanation on how depth adaptation can be achieved with minimal performance degradation. Under this principle, our approach can avoid typical exhaustive training of target sub-network and instead focus on optimizing the sub-paths of the network to have specific properties. At test time, those sub-paths can be connected in a combinatorial manner to construct sub-networks with various accuracy-efficiency trade-offs from a single network. Experimental results show that those sub-networks form a better Pareto frontier than non-adaptive baseline networks and achieve actual inference acceleration.
@@ A.1 Detailed Architectures
Table 4: Each residual stage of base models is evenly divided into two sub-paths; the first is mandatory and the other is skippable. Since ViT does not define hierarchical stages, 12 identical encoder blocks are divided into 4 stages, resembling other residual networks for vision tasks.
@@ A.2 Performance of Sub-Networks
FFFT 3.90 76.2
TTFF 3.46 76.8
TFTF 3.24 75.4
TFFT 3.68 76.1
FTTF 3.02 75.3
FTFT 3.46 75.4
@@
TTTF 2.80 76.1
TTFT 3.24 75.9
TFTT 3.02 75.4
FTTT 2.80 75.6
@@
FFFF 17.58 81.3
TFFF 16.20 80.5
FTFF 16.20 80.2
FFTF 16.20 80.0
FFFT 16.20 80.4
TTFF 14.67 79.9
TFTF 14.67 79.6
TFFT 14.67 79.8
FTTF 14.67 79.4
FTFT 14.67 79.7
FFTT 14.67 79.4
TTTF 13.21 79.3
TTFT 13.21 79.4
TFTT 13.21 79.2
FTTT 13.21 79.1
TTTT 11.76 79.1
@@
Table 5: FLOPs and ImageNet validation accuracy of sub-networks. Only the super-net and the base-net are trained explicitly. The other parameter-sharing sub-networks in the middle can be selected at test time without explicit training. The highest accuracy in each group is shown in bold. This results show that skipping sub-paths in deeper residual stages is more detrimental to the performance in general, implying that refinement of features is more needed for deeper residual stages.
A.3 Varying the Lengths of Sub-Paths in ResNet50-ADN
@@
Table 6: Varying the length ratio of sub-paths of ResNet50-ADN. Total number of blocks of each stage is not changed.
A.4 Varying Sub-Paths Length Ratios in Vit-b/16-ADN
Each stage of Vit-b/16-ADN has 3 encoder blocks and, by default, we select every last encoder block as a skippable sub-path. Therefore every stage has two mandatory blocks and 1 skipable blocks as shown in Table 4-(b). Figure 7-(a) shows different configuration where every last two blocks of the stages become skippable. With this configuration in Figure 7-(a), we can select much smaller sub-networks. For example, the smallest sub-network, or the base-net, of Vit-b/15-ADN has only 4 mandatory blocks and it requires 5.82 GFLOPs. However, the result in Figure 7-(b) shows that this configuration significantly degrades performance of all sub-networks. As demonstrated with ResNet50-ADN in Figure 6, low inference capability of shallow mandatory sub-paths affects all sub-networks. This again shows that some minimum depths in mandatory sub-paths is required for effective inference.
@@ B.2 Visual Analysis of Sub-Paths
(b) ResNet50-ADN (FFFF), or super-net
(c) ResNet50-ADN (TTTT), or base-net
Figure 8: Class Activation Maps of the 3rd residual stages of ResNet50s. (a) Original ResNet50’s activation regions change gradually across all blocks. (b) In ResNet50-ADN (super-net), the first 3 blocks have extensive hot activation regions, implying active learning of new level features. In contrast, the skippable last 3 blocks have far less activation regions and they are gradually refined around the target. (c) Even though parameters are shared, the activation map of base-net is very different from super-net’s since they use different batch normalization operators.

### v3

@@ 4.3 On-Device Performance
Figure 5-(b) shows the performance on Nvidia Jetson Orin Nano. The inference latency and energy consumption of ResNet50-ADN is compared to S-ResNet50, a representative width-adaptation network. The result shows that depth-adaptation of ResNet50-ADN is highly effective in accelerating inference speeds and reducing energy consumption. Although our ResNet50-ADN has a limited FLOPs adaptation range, reducing FLOPs by 38% through depth adaptation reduces both inference latency and energy consumption by 35%. In contrast, even though S-ResNet50 can reduce FLOPs by up to 93% by adjusting its width, it only achieves up to 9% acceleration in practice.
@@ 4.4 Ablation Study
Table 2: Ablation analysis with ResNet50-ADN and ViT-b/32-ADN. Applied components are checked. $\downarrow$ and $\uparrow$ in parentheses are comparisons to non-adaptive individual networks. By default, only the outputs, or $\mathbf{\hat{y}}_{super}$ and $\mathbf{\hat{y}}_{base}$, are used for self-distillation. The last row with double check marks shows the results when both intermediate features and outputs are used for self-distillation.
self- skip-aware ResNet50-ADN Acc@1 (%) ViT-b/32-ADN Acc@1 (%)
@@
76.1% ($\downarrow$ 0.6%) 74.9 % ($\downarrow$ 0.1%) 76.4% ($\uparrow$0.5%) 74.3% ($\uparrow$0.5%)
76.6% ($\downarrow$ 0.1%) 75.1% ($\uparrow$ 0.1%) 76.0% ($\uparrow$0.1%) 74.3% ($\uparrow$0.5%)
77.6% ($\uparrow$ 0.9%) 76.1% ($\uparrow$ 1.1%) 76.6% ($\uparrow$ 0.8%) 74.3% ($\uparrow$0.5%)
✓✓ 77.3% ($\uparrow$ 0.6%) 76.2% ($\uparrow$ 1.2%)
We first investigate the influence of two key components of the proposed adaptive depth networks: (1) self-distillation of sub-paths and (2) skip-aware BNs/LNs. When our self-distillation method is not applied, the loss of the base-net, or $loss_{base}$, in Algorithm 1 is modified to $criterion(\mathbf{y},\mathbf{\hat{y}}_{base})$. Table 2 shows the results. For ResNet50-ADN, when neither of them is applied, the inference accuracy of the super-net and the base-net is significantly lower than non-adaptive individual networks by 1.5% and 2.8%, respectively. This result shows the difficulty of joint training sub-networks for adaptive networks. When one of the two components is applied individually, the performance is still slightly worse than individual networks’. When both self-distillation and skip-aware BNs are applied together, ResNet50-ADN achieves significantly better performance than individual networks, both in the super-net and the base-net. The last row, with double check marks, shows that exploiting intermediate features as well as softmax outputs for self-distillation has minor impact on performance.
@@
Self-Distillation Strategies: Our self-distillation approach in Algorithm 1 exploits only two sub-networks exclusively as a teacher and a student. Specifically, the super-nets, or FFFF, acts as the the teacher and the base-nets, or TTTT, becomes the student. The purpose of this strategy is not to train only those two sub-networks, but rather to train skippable sub-paths in a way that minimally modifies the feature distribution, as demonstrated in Figure 3. To investigate the effect of self-distillation strategies, we conduct an experiment in Table 3. In every training iteration, rather than exclusively using FFFF and TTTT sub-networks for self-distillation, we randomly sample either a teacher, a student, or both from $2^{4}$ sub-networks. These randomly sampled sub-networks are trained explicitly through self-distillation. However, as shown in Table 3, all sub-networks, both from ResNet50-ADN and ViT-b/32-ADN, perform significantly better when our self-distillation strategy is applied for training. Even though most of our sub-networks (such as TFFF, TTFF, and TTTF) are instantly constructed at test time without explicit training, they still outperform their counterpart sub-networks trained explicitly through random sampling. This result demonstrates that our method of training sub-paths is more effective than training target sub-networks themselves.
Figure 6: Pareto frontiers of three ResNet50-ADNs, each trained with varying ratios between mandatory and skippable sub-paths. Total number of blocks remains unchanged.
Lengths of Mandatory Sub-Paths: In Figure 6, we investigate the impact of varying the ratio of lengths between the mandatory and the skippable sub-paths. (Details are in Table 6.) If mandatory sub-paths become shorter than skippable sub-paths, smaller sub-networks can be selected since more layers can be skipped. However, the result in Figure 6 shows that this configuration (shown in blue line) significantly degrades performance of all sub-networks. Since every sub-network shares parameters of mandatory sub-paths, low inference capability of shallow mandatory sub-paths affects all sub-networks. This implies that maintaining certain depths in mandatory sub-paths is crucial for effective inference. Conversely, further increasing the length of mandatory sub-paths (shown in green line) does not further improve performance and only reduces the range of depth adaptation.
@@ 5 Conclusions
We propose a practical approach to adaptive depth networks that can be applied to both CNNs and transformers with minimal training effort. We provide a general principle and a formal explanation on how depth adaptation can be achieved with minimal performance degradation. Under this principle, our approach can avoid typical exhaustive training of target sub-networks and instead focus on optimizing the sub-paths of the network to have specific properties. At test time, these sub-paths can be connected in a combinatorial manner to construct sub-networks with various accuracy-efficiency trade-offs from a single network. Experimental results show that these sub-networks form a better Pareto frontier than non-adaptive baseline networks and achieve actual inference acceleration.
@@ A.1 Detailed Architectures
Table 4: Each stage of base models is evenly divided into two sub-paths; the first is mandatory and the other is skippable. Since ViT does not define hierarchical stages, 12 identical encoder blocks are divided into 4 stages, resembling other residual networks for vision tasks.
@@ A.2 Performance of Sub-Networks
FFFT 3.90 76.7
TTFF 3.46 76.5
TFTF 3.24 75.3
TFFT 3.68 76.4
FTTF 3.02 75.8
FTFT 3.46 75.9
@@
TTTF 2.80 75.9
TTFT 3.24 75.8
TFTT 3.02 75.3
FTTT 2.80 76.0
@@
FFFF 17.58 81.4
TFFF 16.20 81.1
FTFF 16.20 81.0
FFTF 16.20 80.6
FFFT 16.20 81.2
TTFF 14.67 80.9
TFTF 14.67 80.4
TFFT 14.67 80.9
FTTF 14.67 80.5
FTFT 14.67 80.9
FFTT 14.67 80.6
TTTF 13.21 80.5
TTFT 13.21 80.7
TFTT 13.21 80.5
FTTT 13.21 80.6
TTTT 11.76 80.6
@@
Table 5: FLOPs and ImageNet validation accuracy of sub-networks. Only super-net (or, FFFF) and base-net (or, TTTT) are trained explicitly. Sub-networks in the middle can be selected at test time without explicit training. The highest accuracy in each group is shown in bold.
A.3 Varying the Ratio of Sub-Path Lengths in ResNet50-ADN
@@
Table 6: The configurations of ResNet50-ADNs with different proportions between mandatory and skippable sub-paths. Total number of blocks at each stage remains unchanged.
A.4 Varying the Ratio of Sub-Path Lengths in Vit-b/16-ADN
Each stage of Vit-b/16-ADN has 3 encoder blocks and, by default, we select every last encoder block as a skippable sub-path. Therefore every stage has two mandatory blocks and 1 skipable blocks as shown in Table 4-(b). Figure 7-(a) shows different configuration where every last two blocks of the stages become skippable. With this configuration in Figure 7-(a), we can select much smaller sub-networks. For example, the smallest sub-network, or the base-net, of Vit-b/15-ADN has only 4 mandatory blocks and it requires 5.82 GFLOPs. However, the result in Figure 7-(b) shows that this configuration significantly degrades performance of all sub-networks. As demonstrated with ResNet50-ADN in Figure 6, low inference capability of shallow mandatory sub-paths affects all sub-networks. This result again shows that maintaining certain depths in mandatory sub-paths is crucial for effective inference.
@@ B.2 Visual Analysis of Sub-Paths
(b) ResNet50-ADN (FFFF)
(c) ResNet50-ADN (TTTT)
Figure 8: Class Activation Maps of the 3rd residual stages of ResNet50s. (a) Original ResNet50’s activation regions change gradually across all blocks. (b) In ResNet50-ADN (FFFF), the first 3 blocks have extensive hot activation regions, implying active learning of new level features. In contrast, the skippable last 3 blocks have far less activation regions and they are gradually refined around the target. (c) Even though parameters are shared, the activation map of base-net is very different from super-net’s since they use different batch normalization operators.